Write An equation for a square that has an area of 60 with sides of lenght a+b. Use The Distributive property to fully expand the expression
Answer by stanbon(75887)

Write An equation for a square that has an area of 60 with sides of lenght a+b. Use The Distributive property to fully expand the expression
---------
Area = side^2
-----
A = (a+b)^2 = 60
-----
(a+b)^2 = (a+b)(a+b) = a(a+b)+b(a+b) = a^2 + ab + ba + b^2 = a^2 + 2ab + b^2
-----
a^2 + 2ab + b^2 = 60 sq. units
--------------------------------
Cheers,
Stan H.
===============